OpenStudy (mathhelp346):

PLEASE HELP! Graph: 2x+y=1 x-2y=3

OpenStudy (anonymous):

set the equations to y= format so 2x+y=1 would be y=-2x+1, then pick some x values and graph

OpenStudy (mathhelp346):

so it would be: y=-2x+1 y=1/2x-1.5

OpenStudy (anonymous):

ya

OpenStudy (mathhelp346):

ok